### v2

@@ Abstract
Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes. We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow – implementing new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises 5,258 problems from Codeforces and is continuously updated via an automated pipeline, which decomposes each problem into subproblems with unit tests based on dependency tree analysis and dataflow analysis. We further propose a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees. Extensive experiments on 16 popular LLMs reveal significant performance degradation in multi-turn scenarios. For instance, o1-mini retains only 20.8% Pass@1 in multi-turn scenario versus 37.8% in single-turn scenario. More fine-grained analysis illustrates that model performance inversely correlates with dependency complexity. These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval [5] and MBPP [3] establishing foundational standards. As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows [21, 23, 28]. Modern benchmarks such as DevBench [25] and SWE-Bench [22] now emphasize practical scenarios like bug fixing. However, current benchmarks [9, 15, 35] still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
@@
In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development [24, 1, 6]. By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork [12, 7]. For example, React’s core package alone sees over 37 million weekly downloads across 2 300+ dependent modules, illustrating the productivity gains of modular reuse [4]. As shown in Fig. 1 and 2, by building solutions by traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries. For LLMs to integrate effectively into such workflows, they must handle multi-turn interactions, manage dependencies between code segments and generate reusable components for subsequent iterations.
Figure 1: An example of building a currency arbitrage detection and execution system. The left side demonstrates a codeflow workflow, where modular functions are reused to construct the main logic. The right side represents a single-turn implementation, combining all steps into a monolithic block.
@@
Despite the growing demands, current benchmarks have not fully captured the multi-turn and iterative aspects of codeflow. Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation. While recent benchmarks like BigCodeBench [40] and SWE-Bench [22] have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown. Secondly, the few existing multi-turn benchmarks such as MTPB [31], only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies. Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment. Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
To bridge this gap, we propose CodeFlowBench, a novel evaluation benchmark that sources its problems from the competitive programming platforms like Codeforces [^1] https://codeforces.com/, featuring high difficulty (competition-level problems), high quality (peer-reviewed editorial solutions), and periodic updates (frequent problem additions), elaborated in Appendix A. CodeFlowBench leverages platforms’ frequent release of new problems to ensure continuous updates for the uncontaminated benchmark. Based on dependency decomposition and online submission check, an automated, lightweight pipeline is further developed to process original problems into sets of multi-turn, iterative subproblems with verified test cases and solutions. Remarkably, CodeFlowBench also introduces a specialized evaluation framework, including structural labels and metrics derived from dependency trees, to comprehensively assess multi-turn performance and provide more insights for model shortcomings.
@@
In this paper, we present the first edition of CodeFlowBench, which comprises 5,258 complex problems fetched from the Codeforces problem archive accumulated until now. To ensure ongoing fairness and freshness, we will periodically perform the automated pipeline and update the benchmark with brand-new, uncontaminated problems. Our main contributions are threefold:
  • Pipeline Innovation: We develop a data curation pipeline based on competition platform. This pipeline decomposes official solutions based on function dependencies to create multi-turn, iterative code problems that require code reuse. This lightweight pipeline is fully automated, rigorously verifiable, and easy to maintain, facilitating frequent updates.
  • Benchmark Construction: We introduce the first benchmark specifically designed to evaluate the multi-turn, iterative code generation capabilities for codeflow settings. It features a large and diverse collection of 5,258 complex code problems. By leveraging official editorial solutions, automatically verifying generated code, and capturing test cases through input-output streams, we ensure the quantity and quality of our dataset.
  • Evaluation Design & Insights: We propose a novel evaluation framework that contrasts multi-turn and single-turn code generation patterns. Our framework introduces new structural labels and metrics derived from the dependency trees, aiming to better capture the unique characteristics of multi-turn tasks. These innovations bridge the gap between traditional evaluation metrics and the specific challenges posed by iterative code generation. Extensive experiments reveal the huge performance drop in the codeflow scenarios, even for large reasoning models, highlighting the need for more advanced multi-turn capabilities.
@@ 2 Related Work
The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs [26, 32, 39], but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios. Early works like HumanEval [5] and MBPP [3] focus on standalone functions with low complexity and limited dependency environments. Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS [13], LiveCodeBench [20] and SWE-Bench [22] are limited to single-turn code generation or modification. For the few existing multi-turn benchmarks, MTPB [31] focus on overly simplistic single-function programming without paired unit tests, while InterCode [38] discusses interactive coding with execution feedback. In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable. These limitations highlight the crucial gaps for codeflow benchmarking, and the CodeFlowBench pioneers this research line.
@@
Recent years have witnessed unprecedented progress in code generation capabilities of LLMs. Early works such as Codex [5] and AlphaCode [26] demonstrated proficiency in tasks ranging from code completion to competition-level problem solving. With the scaling up of pre-trained models, exemplified by GPT-4 [2], Code-Llama [33], Deepseek-Coder [10] and Qwen2.5-Coder [16], these advanced models have impressive performance across various programming tasks, languages and domains. Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks. Models such as WizardCoder [29] and Magicoder [36] leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder [14] and MapCoder [18] enable autonomous planning, iterative refinement, and self-evaluation. Despite these advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering. Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation [21].
@@ 3 CodeFlowBench
CodeFlowBench is a lively-updated benchmark which currently contains 5,258 diverse, high-quality and challenging problems. An example problem of CodeFlowBench is shown in Figure 2. In this section, we introduce the data curation pipeline and evaluation framework of CodeFlowBench.
Figure 2: An example problem from CodeFlowBench, sourced from Codeforces problem 2055E. The top-left block displays the problem information. The bottom-left block illustrates the dependency structure, with nodes as subproblems and edges as dependency relationships. The right block details each subproblem, where models solve these subproblems iteratively. The bottom block shows the LLM interaction workflow for each turn.
@@ 3.1 Data Curation Pipeline
We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems. The pipeline consists of two main phases: (1) Data Preparation (Stages I–III), which involves source-specific routines tailored to collect and normalize raw problem data from various platforms. (2) Subproblem Generation (Stages IV–VI), which employs an universally applicable problem decompostion framework. This framework systematically analyzes dependency relationships between functions or classes, and back-translates functions into coherent and sequential subproblems. Notably, our implementation is the first to integrate both official problem editorials and automated submission checker into the data curation process. Leveraging official editorials and submission checker provides reliability benefits, such as eliminated copyright issues or data cleaning needs, and it also ensures the codeflow implementation is high-quality and optimal for each problem. As shown in Fig. 3, our pipeline specifically comprises the following stages:
Figure 3: The data curation pipeline. Left: in the data-collection stage, we scrape problems and corresponding text and code solutions. Middle: in the data-processing and subproblem-generation stage, we validate solutions, parse codes into abstract syntax trees, decompose original problems into individual subproblems, and generate test cases. Right: the final labeling stage for complexity analysis, and the curated CodeFlowBench problem set.
Stage I: Scrape Problem Information
We initiate the process by visiting the Codeforces problem list and navigating to each problem’s page to extract the following information: problem ID, title, time and memory limits, problem description, input/output specifications, sample tests, notes. Furthermore, by querying the official Codeforces API, we retrieve each problem’s rating and tags. Please refer to Appendix B.1 for information details and examples.
Stage II: Scrape Editorial Information
We gather editorial content from corresponding tutorial pages, where each problem’s ID, textual explanation, and solution code are listed sequentially. By identifying problem IDs through URL and text patterns, we extract the relevant content and associate it with the problem list generated in Stage I. Low-quality entries are filtered out, and solutions are categorized into text and code components. Please refer to Appendix B.2 for more details.
Stage III: Solution Code Generation
We input both the problem and its explanation information into the Deepseek-R1 model [11], which either directly uses the provided code implementation or organizes the existing textual solution and converts it into parsable solution code that meets the requirements for extracting function dependencies and constructing subproblems. The generated code is then submitted to the Codeforces platform for correctness verification, and we retain only those codes that successfully solve the problem. The prompt template can be found in Appendix B.3.
Stage IV: Subproblem Generation
We parse the abstract syntax tree (AST) of codes from Stage III to extract function dependencies, filter out built-in functions, and conduct topological sort so that lower-level ones come first. We then treat each function as a subproblem and record its solution code, AST depth, and dependencies information. These information, along with the problem and explanation, are fed into Deepseek-V3 [27] to back-translate a subproblem description. Finally, a subproblem contains the following fields: name, statement, depth, dependencies, solution. Detailed explanations of the prompt template and an example can be found in Appendix B.4.
Stage V: Test Cases Generation
For every problem, publicly accessible test cases from the Codeforces platform are scraped and used to execute the respective solution codes. During execution, we capture input-output pairs at every function call location, which serve as test cases for the corresponding subproblems. Further elaboration on this process can be found in Appendix B.5.
Stage VI: Automatic Labeling
To quantify the complexity of the curated problems for fine-grained analysis, we define two metrics: Overall-Turns–the number of turns required to solve the entire problem; Overall-Depth–the depth of the AST, where the main/solve function as the root and subproblems as nodes. As shown in Fig. 4, we report the distributions of overall-turns and overall-depths, and correlate them with the difficulty ratings provided by Codeforces. The strong correlation observed confirms that our label effectively captures problems’ intrinsic complexity.
(a) Distributions of overall-turns and overall-depth.
(b) The Correlations with Rating Levels.
Figure 4: Statistics of the overall-turns and overall-depth metrics in CodeFlowBench. Subfigure (b) shows inflection points at turns = 1 and depth = 1. This is attributed to the fact that competition-level problems are not restricted to multi-turn or deeply nested structures. Partial difficult problems are designed to be solvable by single function such as number theory related problems.
@@ 3.2 Evaluation Framework
The evaluation design of CodeFlowBench follows best practices defined in HumanEval [5] but introduces a novel multi-turn iterative paradigm. That is, models are required to implement a given function in each round. However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models. Beyond the function signature and problem description, we include dependency information and pre-implemented functions for code reuse. Additionally, we provide an overarching background description of the problem. This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally.
As illustrated in Fig. 2, models must implement target functions while leveraging dependency relationships and pre-implemented components from prior turns. Formally, given $T$ total turns, the function implementation $C_{t}$ at turn $t$ is generated based on:
  • Function signature $F_{t}$ and subproblem description $S_{t}$ for the current target function,
  • Background context $B$ describing the overarching programming problem,
  • Dependency specifications $\mathcal{D}=\{k_{1},...,k_{d}\}$ indicating prerequisite functions $F_{k_{1}},...,F_{k_{d}}$,
  • Verified functional implementations $\{C_{1},...,C_{t-1}\}$ of previous dependencies.
@@
$C_{t}\leftarrow p$ $\Bigl{(}\cdot\,\Bigm{|}\,F_{t},S_{t},\{F_{k_{1}},F_{k_{2}},\ldots,F_{k_{d}}\},\{C_{1},C_{2},\ldots,C_{t-1}\},B\Bigr{)}\quad(1\leq t\leq T),$
$\{C_{1},C_{2},\ldots,C_{T}\}\leftarrow p\Bigl{(}\cdot\,\Bigm{|}\,\{F_{1},F_{2},\ldots,F_{T}\},\{S_{1},S_{2},\ldots,S_{T}\},B\Bigr{)}.$
@@
We adopt widely-used Pass@k [5] as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis. While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts - two models may fail solve the entire problem at different stages but receive identical scores.
To address this limitation, we propose a new metric, Pass Turn ($PT$), which identifies the exact turn at which a model fails by leveraging unit tests for each subproblem. However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth. We further define, Pass Depth ($PD=D-d$), for bottom-up programming, where $d$ and $D$ are working depth and total depth. For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth. For multiple trials, we define APD@k and APT@k following the Pass@k expression. Refer to Appendix C for more details.

### v3

@@ Abstract
Modern software development demands code that is maintainable, testable, and scalable by organizing the implementation into modular components with iterative reuse of existing codes. We formalize this iterative, multi-turn paradigm as codeflow and introduce CodeFlowBench, the first benchmark designed to comprehensively evaluate LLMs’ ability to perform codeflow - implementing new functionality by reusing existing functions over multiple turns. CodeFlowBench comprises two complementary components: CodeFlowBench-Comp, a core collection of 5,000+ competitive programming problems from Codeforces updated via an automated pipeline and CodeFlowBench-Repo, which is sourced from GitHub repositories to better reflect real-world scenarios. Furthermore, a novel evaluation framework featured dual assessment protocol and structural metrics derived from dependency trees is introduced. Extensive experiments reveal significant performance degradation in multi-turn codeflow scenarios. Furthermore, our in-depth analysis illustrates that model performance inversely correlates with dependency complexity. These findings not only highlight the critical challenges for supporting real-world workflows, but also establish CodeFlowBench as an essential tool for advancing code generation research.
CodeFlowBench: A Multi-turn, Iterative Benchmark
for Complex Code Generation
Sizhe Wang1,2, Zhengren Wang1,3, Dongsheng Ma1 Yongan Yu1,4, Rui Ling1, Zhiyu Li*3, Feiyu Xiong3, Wentao Zhang*1 1Peking University 2Shanghai University of Finance and Economics 3Institute for Advanced Algorithms Research, Shanghai 4McGill University {lizy,xiongfy}@iaar.ac.cn, wentao.zhang@pku.edu.cn [Uncaptioned image] Dataset    [Uncaptioned image] Code
footnotetext: $\dagger$ Equal contribution; * Corresponding author.
The first author completed this work during an internship at Peking University.
@@ 1 Introduction
Large Language Models (LLMs) have revolutionized code generation, with benchmarks like HumanEval (chen2021evaluating) and MBPP (austin2021program) establishing foundational standards. As LLM capabilities advance, their role in real-world software development has expanded beyond solving toy problems to supporting complex workflows (jiang2024survey; jin2024llms; liu2024large). Modern benchmarks such as SWE-Bench (jimenez2023swe; li2024devbench) now emphasize practical scenarios like bug fixing. However, current benchmarks (gu2024cruxeval; huang2024effibench; wang2025maintaincoder) still overlook the critical aspect of real-world development: the multi-turn and iterative codeflow scenario.
@@
In modern software engineering, multi-turn and iterative workflows are becoming increasingly prevalent, as the cornerstones of best practices like agile development (larman2004agile; abrahamsson2017agile; cram2019agile). By breaking down complex tasks into manageable subproblems, progressively refine solutions and reuse modular functions, developers can achieve faster delivery, reduced redundancy and enhanced maintainability in teamwork (haefliger2008code; feitosa2020code). For example, React’s core package alone sees over 37 million weekly downloads across 2,300+ dependent modules, illustrating the productivity gains of modular reuse boduch2019react. As shown in Fig. 1 and 2, by building solutions via traversing the function dependency tree from the bottom up, the structured codeflow can both enables parallel development and improves readability, testability and maintainability through modular and responsibility boundaries. To integrate effectively into codeflows, LLMs must learn to “look before and after”, namely accurately leveraging pre-implemented functions while ensuring the modularity of codes to generate.
Figure 1: A currency arbitrage example contrasting the modular codeflow workflow (top) with a monolithic single-turn implementation (bottom).
@@
Despite the growing demands, current benchmarks have not fully captured the multi-turn and iterative aspects of codeflow. Firstly, most benchmarks, such as HumanEval and MBPP, focus only on single-turn code generation. While recent benchmarks like BigCodeBench (zhuo2024bigcodebench) and SWE-Bench (jimenez2023swe) have begun to incorporate practical development scenarios, they still remain at the level of single-turn code modification (or two-turn code generation), leaving the multi-turn generation capabilities unknown. Secondly, the few existing multi-turn benchmarks such as MTPB (nijkamp2022codegen), only focus on single-function programming, lacking both unit tests and sufficient complexity to reflect real-world dependencies. Finally, due to the absence of an update mechanism, previous static datasets risk contamination and unreliable assessment. Therefore, there is a pressing need for a more challenging, well-annotated and frequently updated benchmark specially designed for codeflow.
To bridge this gap, we introduce CodeFlowBench, a benchmark suite that integrates two complementary datasets: CodeFlowBench-Comp, which is sourced from competitive programming platforms like Codeforces [^1] https://codeforces.com/ and CodeFlowBench-Repo, which is derived from diverse real-world GitHub repositories via DomainEval (zhu2025domaineval). Utilizing a unified automated pipeline, CodeFlowBench transforms raw data into sequences of multi-turn, iterative subproblems, accompanied by verified solutions and test cases. This dual-component architecture ensures a broad evaluative scope, spanning reasoning-intensive algorithmic challenge and labor-intensive software engineering. Specifically, CodeFlowBench-Comp provides high-difficulty, high-quality problems with periodic updates, while CodeFlowBench-Repo injects real-world complexity into the assessment. To further reveal model deficiencies, CodeFlowBench introduces a specialized evaluation framework and incorporates structural labels and metrics derived from dependency trees, enabling a nuanced analysis of multi-turn performance.
@@
Our contributions are threefold:
  • Pipeline Innovation: We developed a universal data curation pipeline grounded in function dependency analysis. This fully-automated and lightweight framework decomposes monolithic solutions into multi-turn, iterative coding problems that necessitate strategic code reuse. Its design is rigorously verifiable and highly extensible, enabling the transformation of diverse source code into structured codeflow-style challenges.
  • Benchmark Construction: We introduce CodeFlowBench, the first benchmark to evaluate the iterative, multi-turn code generation capabilities of LLMs. By integrating two complementary datasets: CodeFlowBench-Comp (algorithmic depth) and CodeFlowBench-Repo (real-world domain breadth), we provide a comprehensive testbed to examine software engineering proficiency.
  • Evaluation Design & Insights: We propose a novel evaluation framework that directly contrasts multi-turn and single-turn generation patterns. By introducing structural metrics, such as Average Pass Depth (APD) and Dependency Structure Complexity (DSC), derived from dependency trees, our framework captures the unique complexities of iterative tasks. Extensive experiments reveal a significant performance degradation in codeflow scenarios, even among state-of-the-art reasoning models, highlighting a critical frontier for future research and the need for more advanced codeflow programming capabilities.
@@ 2 Related Work
The landscape of code generation benchmarks has evolved from simpler to more complex tasks to keep pace with the rapid development of LLMs (li2022competition; quan2025codeelo; yu2024humaneval), but still fail to comprehensively capture the multi-turn and iterative features of real-world scenarios. Early works like HumanEval (chen2021evaluating) and MBPP (austin2021program) focus on standalone functions with low complexity and limited dependency environments. Recent benchmarks have emerged to evaluate more complex and realistic scenarios, yet have obvious limitations: most benchmarks such as APPS (hendrycks2021measuring), LiveCodeBench (jain2024livecodebench) and SWE-Bench (jimenez2023swe) are limited to single-turn code generation or modification. For the few existing multi‑turn benchmarks, MTPB (nijkamp2022codegen) focus on overly simplistic single-function programming without paired unit tests, while InterCode (yang2023intercode) discusses interactive coding with execution feedback. In stark contrast, codeflow structures the development into multi-turn processes, ensuring each component is maintainable, testable and reusable. These limitations highlight the crucial gaps for codeflow benchmarking, and the CodeFlowBench pioneers this research line.
@@
Recent years have witnessed unprecedented progress in code generation capabilities of LLMs. Early works such as Codex (chen2021evaluating) and AlphaCode (li2022competition) demonstrated proficiency in tasks ranging from code completion to competition-level problem solving. With the scaling up of pre‑trained models, exemplified by GPT (gpt5), Claude (claude4), Deepseek‑Coder (guo2024deepseek) and Qwen‑Coder (yang2025qwen3), these advanced models have impressive performance across various programming tasks, languages and domains. Building on these foundations, the code generation capabilities have further advanced through instruction tuning and agent frameworks. Models such as WizardCoder (luo2023wizardcoder) and Magicoder (wei2023magicoder) leverage instruction tuning to improve intent alignment and interactive dialogue capabilities, while agent frameworks like AgentCoder (huang2023agentcoder) and MapCoder (islam2024mapcoder) enable autonomous planning, iterative refinement, and self-evaluation. Despite advancements, the community still remains unknown about "how well and how deeply LLMs can perform codeflow"—a critical paradigm for real-world software engineering. Our CodeFlowBench thus provides a principled framework for advancing both model development and evaluation.
@@ 3 CodeFlowBench
In this section, we introduce CodeFlowBench’s data curation pipeline and evaluation framework. Fig. 2 provides an example of CodeFlowBench.
Figure 2: An illustrative example from CodeFlowBench (Source: Codeforces 2055E). The figure illustrates the problem definition with its dependency structure (top), the iterative subproblem decomposition (middle), and the corresponding LLM interaction workflow (bottom).
@@ 3.1 Data Curation Pipeline
We designed an automated, lightweight data-curation pipeline to generate complex multi-turn coding problems. As shown in Fig. 3, the pipeline mainly consists a data preparation phase and a subproblem generation phase.
Figure 3: The data curation pipeline of CodeFlowBench. In Data Preparation Phase (Steps I–III), we scrape problem statements and editorial information, followed by generating and validating the canonical solution codes. In Subproblem Generation Phase (Steps IV–VI), we decompose solutions into subproblems via AST parsing, generate corresponding test cases, and perform automatic complexity labeling to achieve the final CodeFlowBench.
Data Preparation Phase
To address the lack of reasoning-intensive programming datasets tailored for codeflow scenarios, this phase collects programming challenges from Codeforces to construct CodeFlowBench-Comp. We first scrape problem metadata and official editorial content from the platform. To ensure data consistency, raw editorials are processed using DeepSeek-R1 (guo2025deepseek) to generate modular, parsable solution code. These generated solutions undergo rigorous verification via the Codeforces submission system; only implementations that pass all official test cases are retained. By anchoring our dataset in official editorials and validating them against a live judge, we guarantee high-quality, correct codeflow implementations for every problem. Further specifics are provided in Appendix A.1A.3.
Subproblem Generation Phase
This phase functions as a general framework for transforming monolithic solutions into the codeflow format. First, we parse the AST of the verified solution to extract and topologically sort function dependencies. We treat each function as a distinct subproblem, employing Deepseek-V3 (liu2024deepseek) to back-translate its logic into a natural language statement. Next, we execute the full solution against public test cases, instrumenting function calls to capture input-output pairs as ground-truth unit tests for each subproblem. Finally, we annotate problem complexity such as dependency structure complexity, metrics that correlate strongly with official difficulty ratings. Implementation details are provided in Appendix A.4A.5.
Remark. To ensure our pipeline is not biased by the LLMs used for subproblem description and code generation, we conducted several ablation studies. We assessed description consistency across models and validated quality via Human-LLM cross-validation. Furthermore, we verified the structural stability of code decomposition across different LLMs. Ablation details are found in Appendix A.6, and cost estimates for pipeline running are in Appendix A.7 for interested readers.
3.2 Benchmark Details
CodeFlowBench-Comp
This component comprises 5,258 competitive programming problems designed to test algorithmic reasoning depth. It emphasizes multi-step logic with up to $\geq$7 interaction turns. Similarly, dependency structures are non-trivial, concentrated at depth 2, indicating that nested function calls are the norm rather than the exception. To validate the difficulty of these structures, we correlated our metrics with official Codeforces ratings. As illustrated in Fig. 12 (Appendix B for space reason), there is a strong positive correlation: as interaction turns and dependency depths increase, the average problem rating rises sharply, approaching 2,900 (Grandmaster level) for tasks with $\geq$7 turns. These statistics confirm that our structural metrics effectively capture the intrinsic complexity of logical problems.
CodeFlowBench-Repo
Complementing the algorithmic focus, CodeFlowBench-Repo evaluates software engineering proficiency in real-world contexts. It is constructed by applying our Subproblem Generation pipeline to DomainEval (zhu2025domaineval), utilizing high-quality GitHub repositories (>>100 stars) as the source. Our curated dataset spans five specialized domains: Basic, Cryptography, Network, System, and Visualization. Through our pipeline, we extended these original contexts into complex codeflow scenarios that reach up to 9 interaction turns and a dependency depth of 4. The final distribution highlights significant structural complexity: 56.6% of problems require multi-turn iterative generation, forcing models to maintain reasoning continuity, while 54.3% involve deep dependency chains, requiring the model to correctly orchestrate hierarchical function calls. Detailed construction protocols and statistics are provided in Appendix C.1, C.2 and C.3.
@@ 3.3 Evaluation Framework
Our evaluation follows best practices defined in HumanEval (chen2021evaluating), where models are required to implement a given function in each round. However, CodeFlowBench introduces distinct differences in the supporting materials provided to the models. We include the function signature $F_{t}$, problem description $S_{t}$ and pre-implemented functions for code reuse. Additionally, we provide an overarching background description of the problem as $B$. This reflects the real-world scenario where developers often possess overall understanding of the entire problem when implementing code incrementally. As illustrated in Fig. 2, models must implement target functions, $C_{t}$ at turn $t$, by leveraging pre-implemented components from prior turns $\{C_{1},...,C_{t-1}\}$.
@@
$C_{1:T}$ $\leftarrow p\Bigl(\cdot\mid F_{1:T},S_{1:T},B\Bigr).$
$C_{t}$ $\leftarrow p\Bigl(\cdot\mid F_{t},S_{t},C_{<t},B\Bigr),\quad 1\leq t\leq T.$
@@
We adopt widely-used Pass@k (chen2021evaluating) as the main metric for both multi-turn and single-turn cases, but augment it with novel diagnostics for multi-turn analysis. While Pass@k evaluates final success rates, the coarse-grained nature fails to capture partial progress in failed attempts -two models may fail solve the entire problem at different stages but receive identical scores.
To address this limitation, we propose a new metric, Pass Turn ($PT$), which identifies the exact turn at which a model fails by leveraging unit tests for each subproblem. However, only considering failing turns may be biased due to the arbitrary topological ordering at the same depth. We further define, Pass Depth (PD=DdPD=D-d), for bottom-up programming, where $d$ and $D$ are working depth and total depth. For statistical significance, we define the Average Pass Turn (APT) and Average Pass Depth (APD) by averaging across problems grouped by turn or depth. For multiple trials, we define APD@k and APT@k following the Pass@k expression. These metrics clearly distinguishing a model’s strong performance on shallow steps from its potential struggles with deep dependencies. Please refer to Appendix D for more details.